Sixx AM Manufacturing has a target debt?equity ratio of 0.52. Its cost of equity is 20 percent, and its cost of debt is 10 percent. If the tax rate is 32 percent, the company's WACC is ______ percent. ( Round your answer to 2 decimal places. )
Answer #1

D/E = 0.52 --> D=0.52E
Total Value, V = D+E = 1.52E
E/V = 1/1.52 = 0.658 , D/V = 1-E/V = 0.342

WACC = E/V*cost of equity + D/V*cost of debt*(1-tax)
WACC = 0.658*20+0.342*10*(1-0.32)= 15.49%